OpenStudy (anonymous):

Over the last century, the population of Springfield has increased by 140%. If the population was 168,000 a century ago, which number is closest to the population today? Select one: a. 385,200 b. 235,200 c. 403,200 d. 168,140

OpenStudy (texaschic101):

168,000 + 1.4(168,000) = your answer

OpenStudy (anonymous):

i got 252000 @texaschic101

OpenStudy (texaschic101):

add that to 168,000

OpenStudy (anonymous):

420000

OpenStudy (texaschic101):

168,000 + 1.4(168,000) = 168,000 + 235200 = 403,200

OpenStudy (anonymous):

thanksssss'

OpenStudy (texaschic101):

no problem :)